Consider Higgins Production which has the following information about its capital structures:

Debt - 1,500, 5 percent coupon bonds outstanding, $1,000 par value, 7 years to maturity, selling for 80 percent of par, the bonds make semi-annual payments

 Common Stock - 100,000 shares outstanding, selling for $45 per share; the beta is 0.80

 Preferred Stock - 25,000 shares of 6 percent preferred stock outstanding, currently selling for $150 per share

 Market Information - 6 percent market risk premium and 4 percent risk-free rate. Required:

Calculate the following if the company has a tax rate of 36 percent

. i. Total Market Value for the Firm

ii. After-tax cost of Debt

iii. Cost of Equity

iv. Cost of Preferred Stock

v. Weighted Average Cost of Capital


1
Expert's answer
2021-04-19T08:15:10-0400

i. Total Market Value for the Firm is: 45,000×100,000 + 150×25,000 = 8,250,000.

ii. After-tax cost of Debt is: 0.05×(1 - 0.36)

iii. Cost of Equity is:

0.04 + 0.8×(0.06 - 0.04)

iv. Cost of Preferred Stock is 0.05.

v. "WACC = (E\/(E + D)\u00d7Re) + ((D\/(E + D)\u00d7Rd) \u00d7(1 \u2013 T)) + P\/(E + D)\u00d7Rp = (4,500,000\/(4,500,000+1,500,000 + 3,750,000))\u00d7(0.04+0.8\u00d7(0.06-0.04))+1,500,000\/9,750,000\u00d70.05\u00d7(1-0.36) + 3,750,000\/9,750,000\u00d70.06=0.0538."